OpenStudy (anonymous):

Rewrite as a single log and simplify ln(x^8)-ln(x^3)

OpenStudy (anonymous):

never mind it would be ln(x^8/x^3) simplify ln(x^5) right?

OpenStudy (anonymous):

awsum, correct

OpenStudy (unklerhaukus):

\[\log a^n=n\log a\]

OpenStudy (anonymous):

yea u can simplify like above after this